### v1

# Neutrinos as a new tool to characterise the Milky Way Centre

Abstract
The Central Molecular Zone (CMZ), a star-forming region rich in molecular clouds located within hundreds of parsecs from the centre of our Galaxy, converts gas into stars less efficient than anticipated. A key challenge in refining star-formation models is the lack of precise mapping of these dense molecular hydrogen clouds, where traditional tracers often yield inconsistent results due to environmental limitations. We demonstrate how, in the not-so-far future, neutrinos will emerge as a robust mass tracer thanks to advancements in neutrino telescopes. Since neutrinos are produced alongside gamma-rays when cosmic-rays interact with molecular clouds, they offer a complementary, systematics-independent measurement of the gas density. In an optimistic case where most gamma-ray emission from the Galactic Centre region originates from pion decays, we expect several tens of muon neutrinos to be detected in about two decades by KM3NeT, Baikal-GVD, and P-ONE combined, which will enable a better determination of the baryonic content in the Galactic Centre region. The CMZ will serve as a testbed to calibrate conventional tracers against neutrinos, ultimately improving gas measurements in distant galaxies, where neutrinos are undetectable, but traditional tracers remain available.
Stars are born in vast, cold clouds of dense gas within galaxies. These clouds, composed primarily of molecular hydrogen, are known as molecular clouds [1]. The efficiency and rate of star formation in molecular clouds are governed by the physical properties of their gas, with density and total mass being the most influential factors [2, 3]. In particular, a correlation between the star-formation rate and cloud surface density has been observed [4, 5, 6, 7] and, to some extent, modelled [2]. However, numerous star-forming clouds, both within and beyond our Galaxy, exhibit anomalous behaviour that is not accounted for by current models [8, 3]. These deviations have been attributed to properties and environmental conditions of the clouds that we are currently unable to measure confidently, such as magnetisation, turbulence, and cosmic-ray heating and ionisation [9, 10]. These factors not only influence star-formation properties but also introduce systematic uncertainties in their measurements [3]. Thus, improving the methods used to reconstruct gas properties and characterise their systematic uncertainties is an essential step towards advancing our understanding of how galaxies convert gas into stars under different environmental and dynamical conditions.
The Central Molecular Zone (CMZ) is the inner 200-pc region at the centre of the Milky Way. This zone contains the densest clouds in our Galaxy and exhibits unique physical conditions, including high temperatures, highly turbulent motions, and strong magnetic fields [8]. Possibly due to these factors, the star-formation rate in the CMZ is lower than model predictions based on its high gas surface density. The CMZ provides a unique opportunity to test and improve the star-formation models, as its relatively close proximity to Earth enables high-precision observations that is not possible for distant galaxies. It also allows the detection of weak signals that remain beyond the sensitivity limits for extragalactic sources. Among these weak signals are gamma-rays and neutrinos, which are produced when cosmic-rays interact with the gas in molecular clouds.
In this article, we explore for the first time how neutrinos from the CMZ can offer new insights into the gas distribution by complementing more traditional probes and gamma-ray observations. These multi-messenger observations have the potential to characterise the biases and uncertainties of these traditional probes, providing a calibration method to improve the accuracy of gas distribution measurements in distant galaxies.
Traditional methods for determining the gas density in molecular clouds rely on indirect probes known as mass tracers. Unlike molecular hydrogen, which is the dominant constituent of the cloud, tracers are species that exist in trace abundances but emit brightly in thermal or line emission. Observing this emission enables the gas density to be estimated through a conversion factor. However, no single tracer is universally optimal, as each has its own characteristics and limitations. For example, carbon monoxide (CO) and dust are commonly used probes because they are abundant and emit brightly in molecular clouds, but their emission easily saturates at high densities [3, 11]. In contrast, tracers such as carbon monosulfide (CS) and hydrogen cyanide (HCN) emit brightly only when a cloud’s density exceeds a certain threshold, making them better suited for tracing gas above the “critical” density [12].
Differences between the gas measurements derived from different mass tracers are significant in the CMZ [c.f. 13, 14, 15, 16]. For example, dust emission indicates a very dense region in the CMZ at around Galactic Longitude (known as Sgr B2), with lower densities elsewhere. On the other hand, CS suggests the gas distribution is more uniform, with the density peaking near the Galactic Centre. The reconstructed gas column densities (i.e. the integrated baryonic mass along the line of sight) are shown in the top panel of Figure 1. Reconciling different tracers has been a challenging problem, and supplementary methods are needed to characterise their uncertainties and establish a way to calibrate them.
Further information about the gas distribution in the Galactic Centre can be extracted from gamma-rays and neutrinos generated when high-energy cosmic-rays interact with the molecular hydrogen in the cloud. Cosmic-rays injected into the CMZ, regardless of the nature or location of their sources, would diffuse due to the presence of the interstellar magnetic field and fill the CMZ. The clouds will hence act as a fixed target and collectively appear as an extended source of gamma-rays and neutrinos. The gamma-ray or neutrino surface brightness at each specific latitude and longitude of the CMZ is correlated with the column density of gas, convolved with the flux of cosmic-rays passing through it. Thus, the measured gamma-ray or neutrino flux can be directly mapped onto a gas distribution, assuming the cosmic-ray density throughout the CMZ is known and these probes are not absorbed.
Over the last decade, independent observations of TeV gamma-rays from the Galactic Centre have been reported by several telescopes, including HESS [18, 19], MAGIC [20], VERITAS [21], and HAWC [22]. This has led the community to establish the existence of three point-like gamma-ray sources in the CMZ. The first two (HESS J1745290 and G0.9+0.1) are associated with the supermassive black hole at the centre of the Galaxy (Sgr A*) and a supernova remnant, respectively. The nature of the third point-like emission (HESS J1746285) is still unsettled and is possibly related to the pulsar wind nebula candidate G0.130.11 [19].
Gamma-ray telescopes have also detected diffuse emission throughout the CMZ, often referred to Galactic Ridge emission within the gamma-ray community. The gamma-ray flux integrated over this spatially extended emission follows a power-law spectrum and does not display a significant energy cut-off, given the accuracy of the available measurements, which are summarised in Figure 2. The morphology of the gamma-ray image is compatible with expectations for the gas distribution traced by CS emission and a cosmic-ray density profile that decreases as the inverse of the distance from the Galactic Centre [18, 19, 20]. This cosmic-ray profile, combined with the lack of a cut-off in the gamma-ray spectrum, has been interpreted as evidence of a source capable of accelerating charged particles up to at least 1 PeV, located within tens of parsecs from the Galactic Centre [18, 19].
While these gamma-ray results are broadly accepted, little emphasis has been placed so far on the choice of mass tracer used to model the gas distribution. In addition, no attempt to constrain the CMZ gas distribution using gamma-ray observations has been reported in the literature. This may be linked to intrinsic limitations that gamma-rays have in determining these quantities. Firstly, gamma-rays could be attenuated by the interstellar radiation field, which can lead to an underestimation of their flux at the source. Secondly, gamma-rays are not exclusively produced in hadronic processes, such as those resulting from cosmic-rays colliding with molecular hydrogen. Indeed, they can also be produced by leptonic processes such as inverse Compton scattering. If present and not accounted for, leptonic processes would lead to an overestimation of a hadronic gamma-ray component. Thus, using the gamma-ray signal to map the CMZ gas distribution requires strong assumptions regarding the lack of gamma-ray attenuation and leptonic contributions.
Neutrinos, on the other hand, are not absorbed, as they interact only via the weak force, and leptonic processes are not typically associated with their production at high energies. Thus, a measurement of the neutrino flux can be mapped to the gas density with a lower systematic uncertainty than any of the methods discussed so far.
Neutrinos, however, are difficult to detect. Their small cross-section explains why the leading km3-scale neutrino telescope, IceCube, has detected only tens of astrophysical neutrinos in more than a decade of operation at the South Pole. Another limitation is the field of view of IceCube, which does not include the Galactic Centre. IceCube searches requiring high angular resolution are limited to the Northern Sky, as a pure sample of “track events” with angular resolution can be extracted with minimal background contamination only if the Earth acts as a filter for atmospheric muons. As a result, IceCube has only provided an initial measurement of the diffuse flux from the Galactic plane, based primarily on a sample of “cascade events” with an angular resolution of above 10 TeV [23].
Fortunately, neutrino astronomy is now at a turning point. Several neutrino telescopes capable of observing the Southern Sky with high angular resolution — down to a tenth of a degree and below — are being constructed or proposed worldwide. Among them, KM3NeT/ARCA in the Mediterranean Sea [24] and Baikal-GVD in Lake Baikal [25] are in an advanced construction phase and are expected to begin operations within the next few years. Other proposed telescopes include P-ONE [26], located offshore of Vancouver Island, as well as TRIDENT [27] and NEON [28] in the South China Sea, and HUNT in Lake Baikal or the South China Sea [29]. Tentative timelines suggest that P-ONE will come online at the beginning of the next decade, followed by TRIDENT, NEON, and HUNT in the subsequent decade. This rapidly growing network of neutrino telescopes will dramatically enhance our neutrino detection capabilities in the near future and expand the field of view for high-angular-resolution neutrino observations.
To quantify the extent to which future neutrino data will help resolve discrepancies among mass tracers, we consider a baseline scenario where neutrinos are emitted from the CMZ with the same gamma-ray flux and power-law energy spectrum reported by HESS. This assumption appears robust, as all commonly considered hadronic processes result in a gamma-ray and all-flavour neutrino flux within a 20% difference [30]. However, it may lead to an overestimation of the neutrino flux if some of the gamma-ray flux originates from leptonic processes.
As most of the gamma-ray emission comes from the inner part of the CMZ, we assume that the best-fit HESS flux measurement [19], which is integrated over a larger window, is applicable to the integrated flux over the range of Galactic longitude and latitude that we model, as shown in Figure 1. In particular, we take the fit parameters and uncertainties to extrapolate the neutrino energy spectra at higher energies and construct the 1 probability intervals. These intervals are obtained by randomly sampling each fit parameter from a normal distribution centred at its best-fit value, with a standard deviation equal to the reported fit uncertainty. We then use the sampled set of parameters to build a probability distribution of the flux at each energy, with quantiles defining the desired probability intervals. The results, shown in Figure 2, illustrate how this construction encompasses the gamma-ray measurements performed by different telescopes across various energy ranges.
As this analysis focuses on muon track events due only to muon neutrinos, we scale-down the gamma-ray flux by a factor of 3 to account for flavour mixing due to neutrino oscillations [31]. We then convolve the resulting fluxes with the detector response for an illustrative neutrino telescope located in the Northern Hemisphere at the site of KM3NeT, obtaining the expected reconstructed energy spectra. The detector response as a function of energy and neutrino arrival direction (i.e. the effective area) is assumed to be the same as that of IceCube. To account for Earth’s rotation, we perform this calculation for different times of the day, removing the periods when the Galactic Centre is outside the detector’s field of view, and then summing the results. We apply the same approach to model the expected background contributions from diffuse astrophysical neutrino and atmospheric neutrino contamination, as measured by IceCube. These calculations follow the methodology and inputs detailed in Ref. [32], and the resulting reconstructed energies are shown in Figure 3.
To keep the discussion general, we present our sensitivity studies for the future network of telescopes in terms of IceCube-equivalent-years, i.e. the effective exposure that a single detector with the efficiency of IceCube would need to achieve the combined sensitivity of the network. KM3NeT, Baikal-GVD, and P-ONE are assumed to have the same detection efficiency as IceCube, while TRIDENT, NEON, and HUNT, being larger, are assumed to correspond to 7.5, 10, and 30 IceCube-equivalent detectors, respectively.
Under these assumptions, we estimate a detection rate of CMZ muon neutrinos with energy above 100 GeV of events per IceCube-equivalent-year for the HESS best-fit flux value. The predicted number of events ranges from 0.5 to 1.3 when considering the fit range. The background contribution is expected to be 2.8 events per IceCube-equivalent-year, which is significant but can be mitigated due to its distinct energy and angular distribution. In particular, above 10 TeV, most of the events are expected to originate from CMZ neutrinos, as shown in Figure 3. Additionally, while background events are expected to have isotropic arrival directions, CMZ neutrinos will exhibit a distinct spatial distribution, such as those shown in the middle panels of Figure 1, which convolve the gas column density with a cosmic-ray density decreasing as the inverse of the distance from Sgr A*.
These estimates translate to an expected rate of 40 muon neutrino events for 50 IceCube-equivalent-years—an exposure that KM3NeT, Baikal-GVD, and P-ONE collectively aim to achieve within the next 20 years. Assuming the larger telescopes come online in 2040, we anticipate collecting over 300 IceCube-equivalent-years of data within the same timeframe, resulting in several hundred muon neutrino events from the CMZ. This would mark a transformative moment for neutrino astronomy, ushering in a new era of precision measurements. An illustration of how a future neutrino image of our Galactic Centre might look is shown in the bottom panels of Figure 1, depicting the spatial distribution of 100 muon neutrinos sampled from the gas density distribution inferred from the CS and dust tracers.
While Figure 1 demonstrates how a 2D neutrino map could provide insights into the gas distribution, the relatively small neutrino sample size raises questions about the practical significance of this information. To evaluate its impact, we conduct an illustrative statistical analysis that addresses the question: What is the significance with which we will be able to discriminate between two gas distribution models?
We frame this statistical problem as a hypothesis test, where the null hypothesis assumes that the gas distribution in the CMZ follows that inferred from CS tracers, while the alternative hypothesis assumes it follows the dust tracer. These two illustrative tracers were selected because they are widely used and often yield the most divergent gas distributions, defining an envelope within which other tracers fall [14, 19]. Accordingly, the spatial distributions considered for CMZ neutrinos in our analysis are those shown in the middle panels of Figure 1, while the atmospheric and diffuse astrophysical neutrino backgrounds are assumed to be uniform and integrated over the same longitude and latitude range. Likewise, the reconstructed energy fluxes are taken from those discussed earlier and shown in Figure 3.
We assess the model discrimination significance using a likelihood-ratio test statistic based on an unbinned likelihood. The likelihood itself is constructed by multiplying the probability of each event by a three-dimensional probability distribution of energy, longitude and latitude, given by the sum of signal (i.e. CMZ neutrinos) and background components (i.e. atmospheric and diffuse astrophysical neutrinos). The signal and background rates used to scale the probability distributions are treated as nuisance parameters and varied in the test statistics to maximise the likelihood of each model. The three-dimensional probability distributions are constructed using the distributions described in the previous paragraph, with energy considered to be independent of latitude and longitude. To account for the presence of gamma-ray point-like emission likely unrelated to molecular clouds, we exclude from the latitude-longitude plane the circular regions where they are located.
Given the small sample size, we employ a parametric bootstrapping technique to generate the probability distribution of the test statistic under each hypothesis. To do so, we generate datasets under both hypotheses by sampling signal and background events from the same three-dimensional probability distributions used in the likelihood, and then compute the test on each dataset. Finally, we extract the model-discrimination significance by identifying the quantile of the test statistic’s probability distribution under the alternative hypothesis that corresponds to the median of the null hypothesis distribution. We perform this bootstrapping procedure for a range of IceCube-equivalent-year values to interpolate the evolution of discrimination significance over time.
The results we obtain are shown in Figure 4. The model discrimination significance increases approximately with the square root of the exposure, in line with the expected scaling of statistical uncertainties. Our baseline result, using a projected angular resolution of [33], reaches a discrimination power for the best-fit HESS energy spectrum after about 50 IceCube-equivalent-years. This is an exposure we expect to accumulate within two decades, even if only a fraction of the planned neutrino telescopes are ultimately constructed. The plot also illustrates the impact of angular resolution, which is, along with exposure, the most important detector parameter.
Our forecast significances are illustrative and intended to qualitatively demonstrate the potential of neutrino observations. They can vary significantly depending on the actual neutrino flux and detector performance (see Figure 4). They also depend on the actual distribution of cosmic-rays and gas throughout the CMZ. Nevertheless, they define the general range of sensitivity we can expect to achieve in the future. Over time, the gas distributions reconstructed from each mass tracer will evolve, hopefully reducing biases and uncertainties. Simultaneously, advancements in analysis techniques and hardware performance — particularly improvements in the timing resolution of photosensors — will enhance the angular resolution of future neutrino telescopes and their sensitivity.
The next decade will witness the next wave of neutrino telescopes coming online alongside the Cherenkov Telescope Array Observatory (CTAO). CTAO will enable a step-change in the precision of gamma-ray measurements of the CMZ [34]. In particular, by detecting photons up to 300 TeV, CTAO will provide crucial insights into the presence of a possible cut-off affecting the neutrino spectra above the range explored by HESS and HAWC. At the same time, the next generation of neutrino telescopes will provide constraints on the gamma-ray flux due to hadronic processes, potentially leading to the identification of gamma-ray sources powered by leptonic processes. The reciprocal feedback between these two messengers will drive transformative progress and push the frontiers of multimessenger astronomy.
In conclusion, our work shows how neutrino observations will provide valuable insights into the gas distribution of the CMZ and the fraction of gamma-rays generated through leptonic and hadronic processes in the Galactic Centre. While achieving high-significance model discrimination may require exposure levels that will not be reached for another two decades, neutrino telescopes will ultimately provide a fully independent test of the gas density inferred from mass tracers, offering the most robust constraints in terms of systematic uncertainties. Improved measurements of gas properties will facilitate the development of star-formation models in the CMZ, and this knowledge can be generalised to modeling systems with similar environmental conditions, including galaxies beyond the Milky Way.
Acknowledgements—The authors would like to thank Foteini Oikonomou, Kate Pattle, and Sheng-Jun Lin for useful discussions. This work has been supported by the Cosmoparticle Initiative of UCL. PCWL is supported by a UCL Graduate Research Scholarship and a UCL Overseas Research Scholarship. MA acknowledges support from the Science and Technology Facilities Council, part of the UK Research and Innovation (Grant No. ST/T004169/1). ERO is an international research fellow under the Postdoctoral Fellowship of the Japan Society for the Promotion of Science (JSPS), supported by JSPS KAKENHI Grant Number JP22F22327, and also acknowledges support from the RIKEN Special Postdoctoral Researcher Program for junior scientists.